# Interpret the Internal States of Recommendation Model with Sparse Autoencoders

### v2

@@ Abstract.
Recommendation model interpretation aims to reveal models’ calculation process, enhancing their transparency, interpretability, and trustworthiness by clarifying the relationships between inputs, model activations, and outputs. However, the complex, often opaque nature of deep learning models complicates interpretation, and most existing methods are tailored to specific model architectures, limiting their generalizability across different types of recommendation models. To address these challenges, we propose RecSAE, an automated and generalizable probing framework that interprets Recommenders with Sparse AutoEncoder. It extracts interpretable latents from the internal states of recommendation models and links them to semantic concepts for interpretation. RecSAE does not alter original models during interpretation and also enables targeted de-biasing to models based on interpreted results. Specifically, RecSAE operates in three steps: First, it probes activations before the prediction layer to capture internal representations. Next, the RecSAE module is trained on these activations with a larger latent space and sparsity constraints, making the RecSAE latents more mono-semantic than the original model activations. Thirdly, RecSAE utilizes a language model to construct concept descriptions with confidence scores based on the relationships between latent activations and recommendation outputs. Experiments on three types of models (general, graph-based, and sequential) with three widely used datasets demonstrate the effectiveness and generalization of RecSAE framework. The interpreted concepts are further validated by human experts, showing strong alignment with human perception. Overall, RecSAE serves as a first step in both model-level interpretations to various types of recommenders without affecting their functions and offering the potential for targeted tuning of models. The code and data are available [^1] https://github.com/Alice1998/RecSAE.
Recommendation Model, Interpretation, Sparse Autoencoder.
ccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. RecSAE employs a probing paradigm to extract interpretable latents from recommenders. The examples are interpretations of SASRec on Amazon.
Transparency is essential for recommender systems as it promotes user trust, fairness, and accuracy (Zhang et al., 2014). Research on recommendation transparency is typically classified into two categories: result-level explanations and model-level interpretations (Zhang et al., 2020a). Result-level approaches focus on explaining relationships between inputs and outputs (Kaffes et al., 2021; Gao et al., 2023), providing users with recommendation explanations. Model-level approaches aim to uncover the decision process of the model itself (Tsang et al., 2020; Du et al., 2023), enabling developers to understand its activations and apply targeted de-biasing to improve model behavior.
While model-level interpretations offer notable advantages, interpreting current recommendation models remains challenging due to the opaque nature of neuron activations in deep learning models (Zhang et al., 2021; Montavon et al., 2018). Most existing model-level interpretations rely on specialized model architectures (Wang et al., 2018b; Hada et al., 2021; Xiao et al., 2017; Du et al., 2023; Zhang et al., 2024), limiting their applicability to different types of recommendation models. As shown in Table 1, three studies aim to develop more generalized interpretation methods. Tsang et al. (Tsang et al., 2020) introduced a feature interaction interpretation method for context-based recommendation models. Another distillation-based approach (Zhang et al., 2020b) learns structural knowledge from an explainable model to embedding-based models. However, it requires joint training with the target model, reducing its adaptability to pre-existing models. Du et al. (Du et al., 2023) propose a clustering-based method, learning interpretable clusters with interaction logs and item tags. However, it sacrifices performance and necessitates model ensemble to compensate. Moreover, the above methods cannot apply targeted modifications to model behaviors. To the best of our knowledge, no existing work provides a generalizable framework for interpreting diverse recommendation models without affecting their performance and simultaneously enabling targeted de-biasing of model behavior.
Table 1. Existing studies on explainable recommendations do not provide model-level interpretations across diverse recommender types and support targeted tuning and de-biasing.
Method Representative Studies Model level General Framework Targeted Tuning
@@
Counter- Tan 21 (Tan et al., 2021) $\sqrt{}$
factual Kaffes 21 (Kaffes et al., 2021) $\sqrt{}$
@@
Zhang 20 (Zhang et al., 2020b) $\sqrt{}$ $\sqrt{}\mkern-9.0mu{\smallsetminus}$Jointly trained
Feature Tsang 20 (Tsang et al., 2020) $\sqrt{}$ $\sqrt{}\mkern-9.0mu{\smallsetminus}$Context-based
@@
Motivated by recent advances in employing large sparse autoencoders (SAEs) to interpret transformer-based large language models (Bricken et al., 2023; Templeton et al., 2024; Gao et al., 2024), we utilize SAEs to the interpretation of recommenders. Unlike LLMs, recommendation models exhibit fundamental differences in architecture and behavior, as further discussed in Section 2.2. Various types of recommenders, including general and graph-based models, do not incorporate transformer structures. However, the effectiveness of SAEs for interpretation has been demonstrated primarily within specific structures in transformers.
@@
In this work, we introduce an effective probing paradigm for interpreting various types of recommendation models without altering their functionalities. It also enables target tuning/de-biasing to models. RecSAE operates in three steps: (1) Probe Internal Activations: RecSAE probes the internal activations preceding the recommendation model’s prediction layer to capture its internal representations. (2) Train RecSAE Module: The RecSAE module is trained on these activations with a larger latent space and under sparsity constraints, making RecSAE latents more interpretable than the original model activations. (3) Construct and Verify Semantic Concepts: RecSAE leverages a language model (Llama-3-8b-Instruct (AI@Meta, 2024)) to construct concept descriptions based on input items, latent activations and recommendation model outputs. These concepts are validated by predicting latent activations on new cases and calculating confidence scores based on prediction accuracy. Human experts further verify the constructed concepts and reach an average relevance score of 3.89/4. Additionally, RecSAE enables targeted tuning of the interpreted models by replacing the original activations at the probing position with the reconstructed activations. In summary, our contributions are threefold:
@@
  • We propose the RecSAE framework with a probing paradigm for interpreting various types of recommendation models. It enables model interpretation without affecting the performance of well-trained models.
@@
  • Experiments on three types of recommendation models — general, graph-based, and sequential — across widely used datasets demonstrate the effectiveness and generalizability of the RecSAE framework. These results highlight RecSAE’s capability to provide model-level interpretations and its potential to facilitate targeted de-biasing of models.
@@ 2.1. Interpretable Recommendation
Result-level interpretation approaches, such as explanation distillation (Xu et al., 2023), LLM augmentation (Gao et al., 2023), and reinforcement learning (Wang et al., 2018a), offer generalizability but do not fully reveal the mechanisms of models. In contrast, model-level interpretation methods generate insights into model structures using techniques like path-reasoning (Wang et al., 2018b; Zhao et al., 2020; Xian et al., 2020), textual modeling (Hada et al., 2021; Li et al., 2021, 2023), attention mechanisms (Xiao et al., 2017; Tal et al., 2021), counterfactual methods (Tan et al., 2021), and clustering (Du et al., 2023). However, these methods are often constrained by dependence on specific model architectures, limiting their generalization. Beyond them, three studies offer a certain degree of generalization. Zhang et al. (Zhang et al., 2020b) transferring structural knowledge from a path-based explainable model. However, as this framework requires target interpretation models to be trained jointly with the explainable model, it limits its adaptability to pre-existing models. Tsang et al. (Tsang et al., 2020) propose a feature-based interpretation method, but the approach is tailored to context-based recommendation models utilizing feature interactions. Du et al. (Du et al., 2023) mine taste clusters from user-item interactions and item tags. However, their clustering paradigm impacts model diversity and performance, requiring ensemble learning with at least three models, thereby increasing training costs.
@@
2.2. Neuron Interpretation in Machine Learning
@@
Interpreting deep learning models to understand their decision-making processes is a critical research area (Liang et al., 2021; Jazayeri and Ostojic, 2021; Sajjad et al., 2022; Foote et al., 2023; Fan et al., 2024). The field of neuron interpretation in machine learning has recently advanced through several stages, including interpreting single-layer transformers (Bricken et al., 2023) and analysis of large language models such as GPT-4 and Claude-3 (Gao et al., 2024; Templeton et al., 2024). Specifically, Anthropic has employed dictionary learning (Tošić and Frossard, 2011) to interpret neuron activations in single-layer transformers. Additionally, several studies have analyzed residual stream activations in transformers using sparse autoencoders (Cunningham et al., 2023; Templeton et al., 2024; Gao et al., 2024), aiming to uncover mono-semantic features and improve the interpretability of language models. OpenAI has also leveraged language models to interpret neuron activations at scale (Bills et al., 2023; Gao et al., 2024).
@@
However, recommendation systems exhibit several fundamental differences. Firstly, not all recommendation models are transformer-based. A wide variety of models—such as general, sequential, graph-based, and context-based, content-based models—exist, each with distinct structural characteristics. While sparse autoencoders have demonstrated effectiveness primarily on residual activations within transformer structures (Templeton et al., 2024; Gao et al., 2024), their applicability to other architectures remains underexplored. Additionally, recommendation models are trained on interaction logs to learn collaborative signals, where ID-based training data lack direct correspondence to semantic concepts. Third, the evaluation of interpretability in recommendation systems differs from that in language models, as items are often associated with explicit attributes, and prior work (Du et al., 2023) employed intra- and inter-attribute similarity for assessment. In this paper, we introduce the RecSAE framework, which is designed to interpret various ranges of recommendation models with semantic concepts and provides similarity- and confidence-based scores to quantify the reliability of these interpretations.
Figure 2. RecSAE Framework. It encodes the inner states of recommendation models into large and sparse latent space and constructs corresponding concepts with confidence scores based on the relationship between recommendation input, latent activations, and recommendation outputs. Example recommendation models are LightGCN (He et al., 2020) and SASRec (Kang and McAuley, 2018).
3. RecSAE Framework
@@ 3.1. Overview
Formally, the interpretation task is defined as follows. Given 1) a recommendation model, $M_{rec}$; 2) ID-based interaction logs $L=\{(ID_{u},ID_{i}),...\}$ (content-based interaction log are also applicable within the framework); 3) item descriptions $D=\{d_{i},...\}$, such as titles or/and categories, the goal is to interpret what information is encoded in the model. Specifically, we aim to uncover relationships between model internal representations and readable texts.
Based on this general setting, we propose RecSAE, a probing framework that uses a sparse autoencoder structure to interpret the internal states of recommendation models. By encoding model activations into a large, sparse latent space, RecSAE uncovers textual concepts that shape recommendation outcomes, providing insights into previously opaque models. As illustrated in Figure 2, RecSAE features a probing architecture and an automated interpretation pipeline, generating textual descriptions for each concept and validating these interpretations with confidence scores. The following subsection 3.2 and 3.3 provide detailed descriptions.
@@ 3.2. RecSAE Training
To achieve interpretable and mono-semantic latents, RecSAE encodes the inner states of recommendation models into a large, sparse latent space through unsupervised learning. This is achieved by using an autoencoder with large latent dimensions and applying sparsity constraints such as the Top-k activation mechanism.
@@ 3.2.1. Probing Site
In recommendation models, user representations are typically formed at the final layer before similarity computation with candidate items. This layer captures the user modeling information encoded by the model. Therefore, as shown in Figure 2, we probe at the last layer prior to prediction in this paper. As discussed in Section 6, future work could explore probing at different sites to gain deeper insights. RecSAE is trained to reconstruct the activations at the probing site.

### v3

# Understanding Internal Representations of Recommendation Models with Sparse Autoencoders

@@ Abstract.
*Corresponding Authors
Recommendation model interpretation aims to reveal the relationships between inputs, model internal representations and outputs to enhance the transparency, interpretability, and trustworthiness of recommendation systems. However, the inherent complexity and opacity of deep learning models pose challenges for model-level interpretation. Moreover, most existing methods for interpreting recommendation models are tailored to specific architectures or model types, limiting their generalizability across different types of recommenders.
In this paper, we propose RecSAE, a generalizable probing framework that interprets Recommendation models with Sparse AutoEncoders. The framework extracts interpretable latents from the internal representations of recommendation models, and links them to semantic concepts for interpretations. It does not alter original models during interpretations and also enables targeted tuning to models. Experiments on three types of recommendation models (general, graph-based, sequential) with four widely used public datasets demonstrate the effectiveness and generalization of RecSAE framework. The interpreted concepts are further validated by human experts, showing strong alignment with human perception. Overall, RecSAE serves as a novel step in both model-level interpretations to various types of recommendation models without affecting their functions and offering potential for targeted tuning of models. The code and data are available [^1] https://github.com/Alice1998/RecSAE.
Recommendation Model, Sparse Autoencoder, Model Interpretation
submissionid: xxx-xxx-xxxccs: Information systems Recommender systems
@@ 1. Introduction
Transparency is critical in recommendation systems, as it fosters trust, fairness, and accuracy (Zhang et al., 2014; Du et al., 2024; Sinha and Swearingen, 2002; Ge et al., 2024). Research on recommendation transparency is typically classified into two categories: result-level explanations (Nilashi et al., 2016) and model-level interpretations (Zhang et al., 2020a). Result-level explanations aim to provide end users with reasons for recommendations (Kaffes et al., 2021; Gao et al., 2023), by revealing relationships between input information and output items. Model-level interpretations aim to uncover the internal decision‐making rationale of recommendation models (Tsang et al., 2020; Du et al., 2023), enabling system designers to understand learned representations and further apply targeted tuning to improve model behavior.
Table 1. Existing studies on recommendation transparency do not provide model-level interpretations across diverse recommender types, and support targeted tuning of models.
Method Representative Studies Model-level General Framework Targeted Tuning
@@
Counterfactual Tan 21 (Tan et al., 2021) $\sqrt{}$
Kaffes 21 (Kaffes et al., 2021) $\sqrt{}$
@@
Zhang 20 (Zhang et al., 2020b) $\sqrt{}$ $\sqrt{}\mkern-9.0mu{\smallsetminus}$ *Jointly trained
Feature Tsang 20 (Tsang et al., 2020) $\sqrt{}$ $\sqrt{}\mkern-9.0mu{\smallsetminus}$ *Context-based
@@
While model-level interpretations offer notable advantages, interpreting current recommendation models remains challenging due to the opaque nature of neuron activations in deep learning models (Zhang et al., 2021; Montavon et al., 2018) and the diversity of recommendation model types (such as general, graph-based, context-based, sequential models) (Patel et al., 2023). Most existing model-level interpretations rely on specialized model architectures or types (Wang et al., 2018b; Hada et al., 2021; Xiao et al., 2017; Du et al., 2023; Zhang et al., 2024; Balog et al., 2019), limiting their applicability to different types of recommendation models. As shown in Table 1, three studies aim to develop more generalized interpretation methods. Tsang et al. (Tsang et al., 2020) introduced a feature interaction interpretation method for context-based recommendation models. Another distillation-based approach (Zhang et al., 2020b) learns structural knowledge from an explainable model to embedding-based models. However, it requires joint training with the target model, reducing its adaptability to pre-existing models. Du et al. (Du et al., 2023) propose a clustering-based method, learning interpretable clusters with interaction logs and item tags. However, it sacrifices performance and necessitates model ensemble to compensate. Moreover, the above methods cannot apply targeted modifications to model behaviors. To the best of our knowledge, no existing work establishes the capability for interpreting diverse types of recommendation models without affecting their performance and simultaneously enabling targeted tuning to models.
Motivated by recent advances in employing sparse autoencoders (SAEs) to interpret transformer-based large language models (Bricken et al., 2023; Templeton et al., 2024; Gao et al., 2024), we utilize the SAE structures to the interpretation of recommenders. SAEs employ large-dimensional, sparsely activated latents to decompose poly-semantic model representations into mono-semantic latents through unsupervised learning objectives, thereby extracting interpretable concepts embedded in the original model. Besides, sparse autoencoders are suitable for recommendation tasks. User–item interactions in recommender systems are inherently sparse, as most items remain unseen by a user. Meanwhile, user interests and profiles are concept- or feature-level rather than item-level (Godoy and Amandi, 2006). While item-level preferences are explicitly observable through user interactions (e.g., clicks or purchases), the underlying concept- or feature-level preferences are latent and not directly reflected in user logs. Therefore, if the model can effectively extract core concept-level representations from sparse item-level data, it would enhance both the accuracy and interpretability of user modeling, as well as the accuracy and interpretability of the resulting recommendations.
Beyond the potential of SAEs in recommendation scenarios, it is worth noting that existing SAE interpretation pipelines have primarily been validated in the context of large language models. However, recommendation models differ fundamentally from LLMs in terms of architecture, training data, and optimization objectives, as further discussed in Section 2.2. Various types of recommenders, including general and graph-based models, do not incorporate transformer structures. However, the effectiveness of SAEs for interpretation has been demonstrated primarily within specific structures in transformers. Consequently, unlocking the potential of SAEs for recommendation model interpretations demands design adaptations and rigorous validation across diverse types of recommendation models.
@@
In this work, we introduce an effective probing paradigm for interpreting various types of recommendation models without altering their functionalities. It also enables target tuning to models. RecSAE interpretation operates in three steps: (1) Probe Internal Representations: RecSAE first probes the internal user representations in recommendation models. (2) Train RecSAE Module: The RecSAE module is trained on these representations with a larger latent space and under sparsity constraints, making RecSAE latents more interpretable than the original model representations. (3) Construct and Verify Semantic Concepts: RecSAE constructs concept descriptions using three sources through two independent methods. The three sources are input items, RecSAE latent activations, and recommendation‐model outputs. The two methods are TF–IDF algorithm (term frequency–inverse document frequency) to derive term‐level interpretations, and a large language model (Llama-3-8b-Instruct (AI@Meta, 2024)) to generate semantic‐level explanations. These interpreted concepts are validated by predicting latent activations on new cases and calculating confidence scores based on prediction accuracy. Human experts further verify the constructed concepts and reach an average relevance score of 3.89/4, showing the efficacy of the automated concept construction pipeline. Additionally, RecSAE enables targeted tuning of the interpreted models by replacing the original representations at the probing position with the reconstructed representations. In summary, our contributions are threefold:
@@
  • We propose the RecSAE framework with a probing paradigm for interpreting various types of recommendation models. This pipeline enables model interpretation without affecting the performance of well-trained models.
@@
  • Experiments on three types of recommendation models — general, graph-based, and sequential — across widely used datasets demonstrate the effectiveness and generalizability of the RecSAE framework. These results highlight RecSAE’s capability to provide model-level interpretations and its potential to facilitate targeted tuning of models.
@@ 2.1. Explainable and Interpretable Recommendation
Result-level interpretation approaches, such as explanation distillation (Xu et al., 2023), LLM augmentation (Gao et al., 2023), and reinforcement learning (Wang et al., 2018a), offer generalizability but do not fully reveal the mechanisms of models. In contrast, model-level interpretation methods generate insights into model structures using techniques like path-reasoning (Wang et al., 2018b; Zhao et al., 2020; Xian et al., 2020), textual modeling (Hada et al., 2021; Li et al., 2021, 2023), attention mechanisms (Xiao et al., 2017; Tal et al., 2021), counterfactual methods (Tan et al., 2021), and clustering (Du et al., 2023). However, these methods are often constrained by dependence on specific model architectures, limiting their generalization. Beyond them, three studies offer a certain degree of generalization. Zhang et al. (Zhang et al., 2020b) transferring structural knowledge from a path-based explainable model. However, as this framework requires target interpretation models to be trained jointly with the explainable model, it limits its adaptability to pre-existing models. Tsang et al. (Tsang et al., 2020) propose a feature-based interpretation method, but the approach is tailored to context-based recommendation models utilizing feature interactions. Du et al. (Du et al., 2023) mine taste clusters from user-item interactions and item tags. However, their clustering paradigm impacts model diversity and performance, requiring ensemble learning with at least three models, thereby increasing training costs. With the development of language-model-based generative recommendation, studies have explored learning semantic IDs to represent items in a discrete, language-like form (Rajput et al., 2023; Wang et al., 2024). These works reveal a misalignment between semantic identifiers and collaborative signals. In contrast, our work focuses on interpreting traditional recommendation models that are purely trained on collaborative information, aiming to understand the semantics implicitly encoded in their latent representations.
@@
2.2. Sparse Autoencoder
@@
Interpreting deep learning models to understand their decision-making processes is a critical research area (Liang et al., 2021; Jazayeri and Ostojic, 2021; Sajjad et al., 2022; Foote et al., 2023; Fan et al., 2024). The field of neuron interpretation has recently advanced through several related work, including interpreting single-layer transformers (Bricken et al., 2023) and analysis of large language models such as GPT-4 and Claude-3 (Gao et al., 2024; Templeton et al., 2024). Specifically, Anthropic has employed dictionary learning (Tošić and Frossard, 2011) to interpret neuron activations in single-layer transformers. Additionally, several studies have analyzed residual stream activations in transformers using sparse autoencoders (Cunningham et al., 2023; Templeton et al., 2024; Gao et al., 2024), aiming to uncover mono-semantic features and improve the interpretability of language models. OpenAI has also leveraged language models to interpret neuron activations at scale (Bills et al., 2023; Gao et al., 2024). Though sparse autoencoders show great potential for interpreting deep learning models, they also face several inherent limitations. In practice, issues such as activation shuffling, over-sparsification, and the lack of guaranteed clean or mono-semantic features remain (Anthropic, 2024; Ann., 2025). Consequently, their effectiveness in interpreting diverse recommendation models trained on inherently sparse user–item interaction data remains largely unexplored.
@@
Besides, recommendation systems exhibit several fundamental differences. Firstly, not all recommendation models are transformer-based. A wide variety of models—such as general, sequential, graph-based, and context-based, content-based models—exist, each with distinct structural characteristics. While sparse autoencoders have demonstrated effectiveness primarily on residual activations within transformer structures (Templeton et al., 2024; Gao et al., 2024), their applicability to other architectures remains underexplored. Additionally, recommendation models are trained on interaction logs to learn collaborative signals, where ID-based training data lack direct correspondence to semantic concepts. Third, the evaluation of interpretability in recommendation systems differs from that in language models, as items are often associated with explicit attributes, and prior work (Du et al., 2023) employed intra- and inter-attribute similarity for assessment. Despite these differences, SAEs hold strong potential in recommendation scenarios. Recommendation models operate on high-dimensional and sparse interaction spaces, while user interests are typically expressed at the feature level rather than the item level. By leveraging sparse autoencoder techniques, user–item co-occurrence patterns can be effectively disentangled into interpretable latent factors aligned with human-understandable concepts. While some prior studies have incorporated SAE-based structures for recommendation tasks (e.g., SLIM (Ning and Karypis, 2011) and EASE (Steck, 2019)), RecSAE differs fundamentally in purpose and scope. These earlier methods aim to build standalone models for improving recommendation accuracy, whereas RecSAE functions as a general interpretation framework—designed to probe and explain existing recommendation models through semantically meaningful latent concepts, thereby enabling both interpretability and targeted tuning. In this paper, we introduce the RecSAE framework, which is designed to interpret various ranges of recommendation models with semantic concepts and provides similarity- and confidence-based scores to quantify the reliability of these interpretations.
3. RecSAE
@@ 3.1. Overview
Formally, the interpretation task is defined as follows. Given 1) a recommendation model, $M_{rec}$; 2) ID-based interaction logs $L=\{(ID_{u},ID_{i}),$...\} (content-based interaction log are also applicable within the framework); 3) item descriptions $D=\{d_{i},$...\}, such as item titles or/and categories, the objective is to use unsupervised learning method to extract mono-semantic latents from the model’s internal representations and interpret them into human-understandable concepts. This process aims to reveal the correspondence between the model’s internal representations and interpretable semantic concepts.
Based on this general setting, we propose RecSAE, a probing framework that uses a sparse autoencoder structure to interpret the internal states of recommendation models. As illustrated in the Introduction, sparse autoencoder techniques are well-suited to recommendation scenarios. User–item interactions are inherently sparse, as most items remain unseen by a given user. However, user interests are typically expressed at the feature or concept level rather than the item level. This motivates concept-level modeling to better interpret user preferences. Sparse autoencoder techniques are well-suited to these scenarios. By encoding model representations into a large, sparse latent space, RecSAE uncovers textual concepts that shape recommendation outcomes, providing insights into previously opaque models. As illustrated in Figure 1, RecSAE features a probing architecture and an automated interpretation pipeline, generating descriptions for each latent and validating these interpretations with confidence scores. The following subsection 3.2 and 3.3 provide detailed descriptions.
Figure 1. RecSAE Framework. It encodes the inner representations of recommendation models into large and sparse latent space and constructs corresponding concepts with confidence scores based on the relationship between recommendation input, latent activations, and recommendation outputs. Example recommendation models are LightGCN (He et al., 2020) and SASRec (Kang and McAuley, 2018).
@@ 3.2. RecSAE Training
In this section, we describe the unsupervised learning process of RecSAE, which extracts interpretable and mono-semantic latents from the internal representations of recommendation models. To promote interpretability, RecSAE employs a large latent dimensionality combined with sparsity-controlled activation mechanism (Gao et al., 2024). This design disentangles the original dense, poly-semantic representations into a larger number of sparse, mono-semantic latents.
In the following sections, we introduce RecSAE’s probing site, module architecture, and the associated training loss function.
@@ 3.2.1. Probing Site
User‐interest representations encode the personalized information learned by recommendation models (Ko et al., 2022). These representations are typically formed before the prediction layer for similarity computations with candidate items. Accordingly, as illustrated in Figure 1, we probe at the user representations prior to the prediction layer. Future work could investigate additional probing sites to gain deeper insights. RecSAE is trained to reconstruct the representations extracted at the probing site. Formally,
[…]
where $M_{rec}$ represents the interpreted recommendation model, $u$ denotes the user identifier or interaction history, $i_{pred}$ represents the model’s predicted item, $S_{Internal}$ is the user‐interest representation at the probing site, $M_{recP1}$ and $M_{recP2}$ denotes the two sub‐networks before and after the probing site.